# Time Scales in Many-Body Fast Neutrino Flavor Conversion

### 23 Dec 2023 v1

@@ I Introduction
Neutrinos play a fundamental role in extreme astrophysical environments, such as core collapse supernovae and neutron star binary mergers, where they, in part, determine the dynamical evolution of the explosion and set the ambient conditions for the ensuing nucleosynthesis, e.g., Refs. [1, 2, 3, 4, 5, 6, 7]. Important neutrino-matter processes active in these scenarios proceed through weak charged-current reactions, and can therefore be strongly affected by inhomogeneities in the flavor composition of the neutrino flux. It is well known that neutrinos can experience flavor oscillations in vacuum (for a recent review, see Ref. [8]) but, for neutrinos with ${\rm MeV}$ energies, these occur over length scales that are typically large compared to the size of the supernova core. In these environments, flavor oscillations are modified by charged-current scattering with charged leptons, leading to the MSW effect [9, 10], and neutral-current neutrino-neutrino scattering which can lead to collective flavor oscillations [11, 12, 13, 14, 15, 16, 17, 18, 19] (see also Refs. [20, 21, 22] for reviews) which can dramatically reduce the oscillation length scales and thus possibly have important impact during the explosion [23, 24]. The physics of collective flavor oscillations is typically discussed in terms of mean-field Boltzmann kinetic equations [25, 26, 27] which assume that neutrinos de-correlate between successive collisions among themselves. A better understanding of correlation effects beyond mean-field descriptions has been a topic of interest since the beginning of this research area [28, 29, 30, 31], and has recently seen a resurgence of efforts [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], aided in part by the use of available NISQ-era [45] quantum computers, and by quantum information tools more generally [46, 47, 48, 49, 50] (see also Refs. [51, 52, 53] for recent reviews). A direct connection between the presence of unstable collective modes, whose amplitude grows exponentially at early times, with the presence of Dynamical Phase Transitions (DPT) [36, 38] has recently been shown. This led to a proposal [38] of a unified description of unstable modes generated by asymmetries in the vacuum term, the so called ”slow-modes” in the neutrino literature, with unstable modes generated by asymmetries in the flavor dependent angular distribution of neutrino velocities, the so called ”fast-modes” This identification was later shown to hold rigorously at the mean-field level [54]. In this work, we extend the previous analysis of Ref. [38], that considered simple angular geometries, to the full multi-angle regime.
@@
Neglecting non-forward scattering processes between neutrinos and the external matter and neutrinos among each other, the Hamiltonian describing flavor evolution of neutrinos in dense environments contains three main contributions (see e.g., Ref. [20] for a review): a one-body term responsible for neutrino flavor mixing in vacuum, a one-body term describing forward scattering between neutrinos and a static matter background (which gives rise to the MSW effect) and, finally, a two-body interaction term describing weak neutral-current forward scattering among neutrinos. In the two-flavor approximation, neutrinos can be described in terms of a $SU(2)$ flavor isospin, with the convention |$\ket{\nu_{e}}=\ket{\uparrow}$ for the electron flavor and |$\ket{\nu_{x}}=\ket{\downarrow}$ for the heavy flavor. In terms of flavor isospin, the Hamiltonian for $N$ neutrinos can then be written in the flavor basis as [55, 56],
@@
with $\theta_{i}$ the polar angle of the $i^{\rm th}$ neutrino momentum, and where we introduce, for later convenience, the neutrino-neutrino coupling matrix $\mathcal{J}_{ij}$. In our numerical simulations, we use a forward-peaked uniform angular distribution with $\theta_{i}\in[-0.5,0.5]$. We have verified that the conclusions reached with this choice of distribution remain unchanged qualitatively when adopting a different angular range or other smooth non-uniform distributions (such as triangular or Gaussian). As pointed out in previous works [42, 44], the use of uniform grids of angles can lead to undesirable artificial degeneracies in the Hamiltonian’s energy level. To prevent such (lattice) artifacts, the angles $\theta_{i}$ are randomly distributed uniformly on the interval $[0,0.5]$, and ensemble calculations with different realizations are performed. As shown in Ref. [54], the Hamiltonian in Eq. (2) is integrable, in the sense that it allows for an extensive number of conserved charges, and is therefore not able to thermalize the reduced one-body density matrices to a Boltzmann distribution [44]. Nevertheless, it is a model commonly employed in studies of fast-flavor conversion [57, 54], and provides a useful and interesting test case to analyze the time-scales associated to flavor oscillations in the many-body setting.
@@ III Loschmidt Echoes
In the first part of this section, we provide a brief introduction to quantum quenches, the Loschmidt echo, and how Loschmidt echos can characterize DPTs via non-analyticities in Loschmidt rate functions. In the second part, we present a generalization of Loschmidt echo-defined DPTs for our system, and give a brief summary of the computational methods used for the simulations of our neutrino system Eq. (2).
@@ IV Analysis of Loschmidt Echo Crossing Times
where $a_{i}$ denote the real numbers denoting their amplitudes. As defined in Mathematica, the type-1 stable distribution is,
@@
There are some interesting points to note from our results. First is that the locations of the two peaks, $\mu_{1,2}$, become approximately independent of the number of neutrinos, and the variances also appear to be approximately stable. Further, the relative strength of the two distributions appears somewhat stable to increasing neutrino number. The goodness-of-fit to two stable distributions depends on the number of neutrinos, as can be seen in the last column of Table 1. For the systems with $N\geq 8$, there are two peaks in the histograms that are reasonably well reproduced by the sum of two stable distributions. However, as the $\chi^{2}/{\rm dof}>1$ for each, this does not constitute a perfect description of the systems, as expected. It is found to be a superior hypothesis to that of a single stable distribution for $N\geq 8$, which each have $\chi^{2}/{\rm dof}>10$.
It is enlightening to compare the results for the Loschmidt echo crossings with the results of the stability analysis that we presented above, in particular the distributions of imaginary parts of the frequencies. In both instances, there are two peaks in the distributions associated with the sampled ${\cal J}_{ij}$ matrices. The slower and faster modes correspond to the long-time scale and short-time scale Loschmidt echo crossings, respectively. Interestingly, while the distributions associated with faster-modes and the shorter Loschmidt echo crossing times occur at the same $w_{\times}$, those associated with the slower modes (consistent with zero frequency) are at a desperately larger value of $w_{\times}$ than the corresponding longer Loschmidt echo crossing times. The well-separated slower- and faster-mode distributions differ from the overlapping distributions in the Loschmidt echo crossing times. So while the linear instability analysis describes well the fast modes, it is only qualitatively reflective of the longer Loschmidt echo crossing times.
@@ VI Acknowledgements
We thank the following colleagues for their valuable discussions, feedback, and suggestions throughout this work: Joe Carlson, Anthony Ciavarella, Vicenzo Cirigliano, Roland Farrell, Marc Illa, Saurabh Kadam, David Kaplan, Natalie Klco, Joshua D. Martin, Niklas Mueller, Hersh Singh, Jesse Stryker, Francesco Turro, Xiaojun Yao, and the InQubator for Quantum Simulation (IQuS) for providing an intellectually stimulating environment. We would like to thank the QC4HEP Working Group at CERN and the Thrust-2 team in the Quantum Science Center (QSC) (https://qscience.org), a National Quantum Information Science Research Center of the U.S. Department of Energy, that are focused on quantum information and simulation aspects of neutrino physics. This work was supported in part by U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) [^3] https://iqus.uw.edu/ under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science [^4] https://science.osti.gov/np/Research/Quantum-Information-Science This work is also supported, in part, through the Department of Physics [^5] https://phys.washington.edu and the College of Arts and Sciences [^6] https://www.artsci.washington.edu at the University of Washington. We have made extensive use of Wolfram Mathematica [73], python [74, 75], julia [76], jupyter notebooks [77] in the Conda environment [78]. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington. [^7] https://itconnect.uw.edu/research/hpc

### 05 Mar 2024 v2

@@ I Introduction
Neutrinos play a fundamental role in extreme astrophysical environments, such as core collapse supernovae and neutron star binary mergers, where they, in part, determine the dynamical evolution of the explosion and set the ambient conditions for the ensuing nucleosynthesis, e.g., Refs. [1, 2, 3, 4, 5, 6, 7]. Important neutrino-matter processes active in these scenarios proceed through weak charged-current reactions, and can therefore be strongly affected by inhomogeneities in the flavor composition of the neutrino flux. It is well known that neutrinos can experience flavor oscillations in vacuum (for a recent review, see Ref. [8]) but, for neutrinos with ${\rm MeV}$ energies, these occur over length scales that are typically large compared to the size of the supernova core. In these environments, flavor oscillations are modified by charged-current scattering with charged leptons, leading to the MSW effect [9, 10], and neutral-current neutrino-neutrino scattering which can lead to collective flavor oscillations [11, 12, 13, 14, 15, 16, 17, 18, 19] (see also Refs. [20, 21, 22] for reviews) which can dramatically reduce the oscillation length scales and thus possibly have important impact during the explosion [23, 24]. The physics of collective flavor oscillations is typically discussed in terms of mean-field Boltzmann kinetic equations [25, 26, 27] which assume that neutrinos de-correlate between successive collisions among themselves. A better understanding of correlation effects beyond mean-field descriptions has been a topic of interest since the beginning of this research area [28, 29, 30, 31], and has recently seen a resurgence of efforts [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] (see Ref. [45] for a recent review, and also Refs. [46, 47] for recent discussions concerning the many-body approaches) aided in part by the use of available NISQ-era [48] quantum computers and by quantum information tools more generally [49, 50, 51, 52, 53] (see also Refs. [54, 55, 56] for recent reviews). A direct connection between the presence of unstable collective modes, whose amplitude grows exponentially at early times, with the presence of Dynamical Phase Transitions (DPT) [36, 38] has recently been shown. This led to a proposal [38] of a unified description of unstable modes generated by asymmetries in the vacuum term, the so called ”slow-modes” in the neutrino literature, with unstable modes generated by asymmetries in the flavor dependent angular distribution of neutrino velocities, the so called ”fast-modes” This identification was later shown to hold rigorously at the mean-field level [57]. In this work, we extend the previous analysis of Ref. [38], that considered simple angular geometries, to the full multi-angle regime.
@@
Neglecting the non-forward scattering processes between neutrinos and the external matter and neutrinos amongst each other, the Hamiltonian describing flavor evolution of neutrinos in dense environments contains three main contributions (see Ref. [20] for a review): a one-body term responsible for neutrino flavor mixing in vacuum, a one-body term describing forward scattering between neutrinos and a static matter background (which gives rise to the MSW effect) and, finally, a two-body interaction term describing weak neutral-current forward scattering among neutrinos. In the two-flavor approximation, neutrinos can be described in terms of a $SU(2)$ flavor isospin, with the convention |$\ket{\nu_{e}}=\ket{\uparrow}$ for the electron flavor and |$\ket{\nu_{x}}=\ket{\downarrow}$ for the heavy flavor. In terms of flavor isospin, the Hamiltonian for $N$ neutrinos can then be written in the flavor basis as [58, 59],
@@
with $\theta_{i}$ the polar angle of the $i^{\rm th}$ neutrino momentum, and where we introduce, for later convenience, the neutrino-neutrino coupling matrix $\mathcal{J}_{ij}$. In our numerical simulations, we use a forward-peaked uniform angular distribution with $\theta_{i}\in[-0.5,0.5]$. We have verified that the conclusions reached with this choice of distribution remain unchanged qualitatively when adopting a different angular range or other smooth non-uniform distributions (such as triangular or Gaussian). As pointed out in previous works [42, 44], the use of uniform grids of angles can lead to undesirable artificial degeneracies in the Hamiltonian’s energy level. To prevent such (lattice) artifacts, the angles $\theta_{i}$ are randomly distributed uniformly on the interval $[0,0.5]$, and ensemble calculations with different realizations are performed. As shown in Ref. [57], the Hamiltonian in Eq. (2) is integrable, in the sense that it allows for an extensive number of conserved charges, and is therefore not able to thermalize the reduced one-body density matrices to a Boltzmann distribution [44]. Nevertheless, it is a model commonly employed in studies of fast-flavor conversion [60, 57, 61], and provides a useful and interesting test case to analyze the time-scales associated to flavor oscillations in the many-body setting.
@@ III Loschmidt Echoes
In this section, we provide a brief introduction to quantum quenches, the Loschmidt echo, and how Loschmidt echos can characterize DPTs via non-analyticities in Loschmidt rate functions. For a brief summary of the computational methods used for the simulations of our neutrino system Eq. (2) and a generalization of Loschmidt echo-defined DPTs for our system, see section IV.
@@ IV Analysis of Loschmidt Echo Crossing Times
where $a_{i}$ denote their amplitudes and are real numbers. As defined in Mathematica, the type-1 stable distribution is,
@@
There are some interesting points to note from our results. First is that the locations of the two peaks, $\mu_{1,2}$, become approximately independent of the number of neutrinos, and the variances also appear to be approximately stable. Further, the relative strength of the two distributions appears somewhat stable to increasing neutrino number. The goodness-of-fit to two stable distributions depends on the number of neutrinos, as can be seen in the last column of Table 1. For the systems with $N\geq 8$, there are two peaks in the histograms that are reasonably well reproduced by the sum of two stable distributions. However, as the $\chi^{2}/{\rm dof}>1$ for each, this does not constitute a perfect description of the systems, as expected. It is found to be a superior hypothesis to that of a single stable distribution for $N\geq 8$, which have $\chi^{2}/{\rm dof}>10$.
It is enlightening to compare the results for the Loschmidt echo crossings with the results of the stability analysis that we presented above, in particular the distributions of imaginary parts of the frequencies. In both instances, there are two peaks in the distributions associated with the sampled ${\cal J}_{ij}$ matrices. The slower and faster modes correspond to the long-time scale and short-time scale Loschmidt echo crossings, respectively. Interestingly, while the distributions associated with faster-modes and the shorter Loschmidt echo crossing times occur at the same $w_{\times}$, those associated with the slower modes (consistent with zero frequency) are at a desparately larger value of $w_{\times}$ than the corresponding longer Loschmidt echo crossing times. The well-separated slower- and faster-mode distributions differ from the overlapping distributions in the Loschmidt echo crossing times. So while the linear instability analysis describes well the fast modes, it is only qualitatively reflective of the longer Loschmidt echo crossing times.
@@ VI Acknowledgements
We thank the following colleagues for their valuable discussions, feedback, and suggestions throughout this work: Joe Carlson, Anthony Ciavarella, Vicenzo Cirigliano, Roland Farrell, Marc Illa, Luke Johns, Saurabh Kadam, David Kaplan, Natalie Klco, Joshua D. Martin, Niklas Mueller, Hersh Singh, Jesse Stryker, Francesco Turro, Xiaojun Yao, and the InQubator for Quantum Simulation (IQuS) for providing an intellectually stimulating environment. We would like to thank the QC4HEP Working Group at CERN and the Thrust-2 team in the Quantum Science Center (QSC) (https://qscience.org), a National Quantum Information Science Research Center of the U.S. Department of Energy, that are focused on quantum information and simulation aspects of neutrino physics. This work was supported in part by U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) [^3] https://iqus.uw.edu/ under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science [^4] https://science.osti.gov/np/Research/Quantum-Information-Science This work is also supported, in part, through the Department of Physics [^5] https://phys.washington.edu and the College of Arts and Sciences [^6] https://www.artsci.washington.edu at the University of Washington. We have made extensive use of Wolfram Mathematica [77], python [78, 79], julia [80], jupyter notebooks [81] in the Conda environment [82]. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington. [^7] https://itconnect.uw.edu/research/hpc